# sample.bib

### v1

@@ 2.1.3 Generation of Actions
Xu et al.[75] 2023 Vedio, Text Prompt, Control Signal Decision, Control Signal SL Video sequences undergo tokenization leveraging a tokenizer, while text and control signals share a common tokenizer. Following tokenization, the model can generate responses to human inquiries and predict control signals.
$\bullet$ This “SL” refers to “Supervised Learning”.
2.2 End-to-end Autonomous Driving with Pretraining Backbone
As mentioned in 2.1, LLMs-based autonomous driving-related research mainly focuses on improving the interpretability of autonomous driving decisions, enhancing scenario understanding, and guiding the planning or control of autonomous driving systems. The research idea of “pre-training + fine-tuning” in autonomous driving research does not only appear after the introduction of large models but has been researched for a long time. To use a more familiar term, it is end-to-end autonomous driving. In the past few years, scholars have been optimizing Pretraining Backbone in various ways, including the transformer architecture and self-supervised learning methods——Note that we define Pretraining Backbone here as a model that transforms each modal input into a usable feature representation for downstream tasks (such as target detection, trajectory prediction, decision planning, etc.). There have also been many research attempts to develop end-to-end frameworks based on the transformer architecture, with excellent results. Hence, to summarize the application of the underlying models in autonomous driving more comprehensively, we believe that it is necessary to introduce the research related to end-to-end autonomous driving based on Pretraining Backbone. In this section, we review the latest research on Pretraining Backbone with end-to-end autonomous driving solutions, the pipeline of such methods is briefly illustrated in Fig. 4.
Figure 4: The pipeline diagram for the end-to-end autonomous driving system with pretraining backbone. Multimodal sensing information is input to pretraining backbone to extract features, after which it enters into the framework of autonomous driving algorithms built by various methods to realize tasks, such as planning/control, to achieve end-to-end autonomous driving.
2.2.1 Pretraining Backbone
In the end-to-end modeling, feature extraction of low-level information from raw data determines the potential of subsequent model performance to a certain extent, and an excellent Pretraining Backbone can endow the model with more powerful feature learning capability.
Pretraining convolutional networks such as ResNet[79], and VGG[80] are the most widely used backbone for visual feature extraction in end-to-end models. These pretraining networks are often trained leveraging target detection or segmentation as the task to extract generalized feature information, and their competitive performances have been verified in many works. ViT[6] first applied the transformer architecture to image processing and achieved excellent classification results. Transformer has the advantage of optimized algorithms for handling large-scale data with its simpler architecture and faster inference speed. The self-Attention mechanism is very suitable for processing time series data, there is the possibility of modeling and predicting the temporal motion trajectories of objects in the environment, and is conducive to the fusion of heterogeneous data from multiple sources, such as LiDAR point clouds, images, maps, etc.
Another class of Pretraining Backbone, represented by LSS[81], BEVDet[82], BEVformer[83], BEVerse[84], etc., expanded the usability by extracting the images captured by the surround-view camera and converting them to bird’s-eye view (BEV) features through model learning, indexing the local image features from the two-dimensional viewpoint to the three-dimensional space. In recent years, BEV has attracted extensive attention due to its ability to describe the driving scene more accurately, and the research of leveraging BEV features such as Pretraining Backbone output is not limited to camera, and the extraction and fusion of multimodal sensing BEV features represented by BEVfusion[85] has further provided a wider vision for autonomous driving systems. However, it should be pointed out that although the transformer architecture brings great performance enhancement, these Backbone still constructs pre-trained models with supervised learning, relying on massive labeled data, and the quality of the dataset also greatly affects the final result of the model.
In both camera and point cloud processing domains, some works implement pre-trained Backbone by unsupervised or self-supervised learning. Wu et al.[86] proposed the PPGeo model, which uses a large number of unlabeled driving videos to accomplish the pre-training of the visual coder in two stages, and can be adapted to different downstream end-to-end autonomous driving tasks. Sautier et al.[87] proposed BEVContrast for self-supervision of 3D Backbone on automotive LiDAR point clouds, which defines contrasts at the level of 2D Cells in the BEV plane, retaining the simplicity as in PointContrast[88] while maintaining good performance in downstream driving tasks.
2.2.2 End-to-end Autonomous Driving Models
Early work on end-to-end autonomous driving modeling was mainly based on various types of deep neural networks, which were constructed through imitation learning[89, 90, 91, 92] or reinforcement learning[93, 94, 95] methods. The work of Chen et al.[96] analyzed the key challenges facing end-to-end autonomous driving from a methodological perspective, pointing out the future trend of empowering end-to-end autonomous driving with fundamental models such as transformer. Some scholars have tried to build an end-to-end autonomous driving system with transformer and got competitive results. For instance, there have been Transfuser[97, 98], NEAT[99], Scene Transformer[100], PlanT[101], Gatformer[102], FusionAD[103], UniAD[104] and a host of end-to-end frameworks developed based on transformer architecture.
Chitta et al.[97, 98] proposed Transfuser, which takes RGB images and BEV views from LiDAR as inputs, uses multiple transformers to fuse the feature maps, and predicts the trajectory points for the next four steps through a single-layer GRU network, followed by longitudinal and transverse PIDs to control the vehicle operation. NEAT[99] further mapped the locations of BEV scene coordinates to trajectory points and semantics and compresses high-dimensional image features leveraging an intermediate attention map, which allows the model to selectively focus on driving-relevant regions and ignore driving task-irrelevant information. PlanT proposed by Renz et al.[101] used simple object-level representations (vehicles, roads) as inputs to the transformer encoder, and used speed prediction of surrounding vehicles as a secondary task to predict future waypoint trajectories. UniAD proposed by Hu et al.[104] enhanced the design of the decoder and achieving the integration of the full stack of autonomous driving tasks into a unified framework to improve autonomous driving performance, although still relying on different sub-networks for each task. This work also won the CVPR 2023 Best Paper Award, which shows the academic recognition of the end-to-end autonomous driving paradigm.
Moreover, Drive Anywhere proposed by Wang et al.[105] not only realize end-to-end multimodal autonomous driving but also combine with LLM to be able to provide driving decisions based on representations that can be queried through images and texts. Dong et al.[106] generated image-based action commands and explanations by building a feature extraction model based on transformer. Jin et al.[107] proposed the ADAPT model to directly output vehicle control signals with inference language descriptions through an end-to-end model. This is the first driving action captioning architecture based on an action-aware transformer. It accomplished the driving control task while adding natural language narratives to guide the decision-making and action process of the autonomous driving control module. It also helped the user to get the vehicle’s state and the surrounding environment at all times, and to batter understand the basis of the actions taken by the autonomous driving system, which improved the interpretability of the decision-making. This provides a glimpse of the potential of the transformer architecture to enhance the interpretability of end-to-end driving decisions.
3 Data Augmentation
With the further development of deep learning, the performance of FMs with pre-training and fine-tuning as the underlying architecture is getting better and better, and FMs represented by GPT are leading the learning paradigm to shift from rule-driven to data-driven. The significance of data as a key link in model learning is self-evident. In the training and testing process of autonomous driving models, a large amount of scenario data is used to enable the models to have a good understanding and decision-making ability for various road and traffic scenarios. The long-tail problem faced by autonomous driving is also the very fact that there are endless unknown edge scenarios that make the generalization ability of the model seem perpetually insufficient, leading to poor performance. In Table LABEL:tab:2 we provide a brief overview of some representative works.
Data augmentation is crucial for improving the generalization ability of autonomous driving models. The realization of data augmentation needs to consider two aspects: on the one hand, how to obtain large-scale data so that the data fed to the autonomous driving model is diverse and extensive; on the other hand, how to obtain as much high-quality data as possible so that the data used to train and test the autonomous driving model is accurate and reliable. Related research have also roughly chosen two directions to enhance the autonomous driving data, one is to enrich the data content of existing datasets and enhance the data features of driving scenarios, and the other is to generate driving scenarios with multiple levels through simulation. In the following, a review of related research on enhancing data based on language and vision FMs will be presented, in section 3.1 we describe related work on extending datasets, and in 3.2 we describe related work on generating driving scenarios.
3.1 Expansion of Autonomous Driving Datasets
Existing autonomous driving datasets are mostly obtained by recording sensor data and then labeling the data. The features of the data obtained in this are usually low-level and exist more at the level of numerical representation, which is insufficient for the visuo-spatial features of the autonomous driving scenarios. Natural language descriptions are seen as an effective way to enhance scene representation[4], Flickr30k[108], RefCOCO[109], RefCOCOg[110], and CLEVR-Ref[111] use concise natural language descriptions to identify the corresponding visual regions in an image. Talk2Car[112] fused image, radar, and LiDAR data to construct the first object-referenced dataset containing natural language commands for self-driving cars. However, the Talk2Car dataset allowed only one object to be referenced at a time. CityFlow-NL[113] constructed a dataset for multi-target tracking through natural language descriptions, and Refer-KITTI[114] achieved prediction of arbitrary target tracking by leveraging natural language queries in the corresponding task.
FMs represented by language models provide new ideas for enriching and expanding autonomous driving datasets under their advanced semantic understanding, reasoning, and interpretation capabilities. Qian et al.[115] created NuScenes-QA, a visual question and answer dataset for autonomous driving in 3D multi-view driving scenarios, by encoding question descriptions through a language model and obtaining answers through feature fusion with sensor data. Significant progress was made in the use of natural language prompts. Wu et al.[116] extended NuScenes-QA by constructing the dataset NuPrompt by capturing and combining natural language elements and then invoking LLM to generate the descriptions. The dataset provided a finer match between the 3D instances and each of the prompts, which helped to more accurately characterize objects in the autopilot images, and improved the accuracy of the model. In addition to directly extending the augmented autonomous dataset, some scholars have also integrated the chain-of-thinking capability of LLM and the cross-modal capability of the vision model to build an automatic annotation system, OpenAnnotate3D[117], which can be used for multimodal 3D data. Expanding the dataset by utilizing the advanced understanding, reasoning, and interpretation capabilities of the underlying models can help to better assess the interpretability and control of the autonomous driving system, thus improving the safety and reliability of the autonomous driving system.
3.2 Generation of Driving Scenarios
Driving scenarios are of great significance for autonomous driving, and to obtain diverse driving scenario data, real-time acquisition of vehicle sensors alone requires huge costs, and it is difficult to obtain sufficient scenario data for some edge scenarios. Generating realistic driving scenarios through simulation has attracted the attention of many researchers, and traffic simulation research can be mainly divided into rule-based and data-driven. Rule-based approaches[118, 119, 120, 121] require the use of predefined rules, which are generally insufficient for the characterization of complex driving scenarios, and the simulated driving scenarios are simpler and less generalizable. Data-driven approaches[122, 123, 124, 125] refer to the use of driving data to train models, from which the models can continuously learn and adapt. However, data-driven approaches often require a large amount of labeled data for training, which hinders the further development of traffic simulation. The current emergence of some high-quality datasets KITTI[126], APolloScape[127], and nuScenes[128] has somewhat remedied this limitation. However, how to construct an effective learning paradigm to get rid of data labeling becomes more important. Recently, great success has been achieved with generative large models obtained by training based on natural language data, and such models can utilize self-supervised learning to learn a wide range of data.
Enhancing data-driven approaches through generative FMs, such as language and vision models, to accelerate traffic simulation has attracted significant research attention. Marathe et al.[129] efficiently generated a dataset comprising 16 weather extremes by constructing a cueing paradigm leveraging a visual-linguistic model. However, the model had some extension constraints due to the phenomenon of pre-selected fixation in data selection. In response to the fact that some long-tailed scenes can never be collected in multi-view shots, Yang et al.[130] fused verbal cues, BEV sketch, and multi-view noise to design a two-stage generative network BEVControl for synthesizing realistic street scene images. However, BEVControl is insufficient for modeling foreground and background detail information. To address the difficulty of obtaining large-scale BEV representations, Li et al.[131] proposed a spatiotemporally consistent diffusion framework, DrivingDiffusion, to autoregressively generate realistic multiview videos controlled by 3D layouts. The quality of the generated data can be effectively improved by introducing local cue inputs into the vision model. Based on the powerful language understanding, reasoning, and interpretation capabilities of LLM, it has also become a research hotspot to combine LLM into driving scenario generation. Chen et al.[132] realized the combination of numeric vector modality and natural language by pairing control commands collected by reinforcement learning intelligence and question answers generated by LLM to directly construct new data. Zhong et al.[133] proposed a scenario-level diffusion-based language-guided traffic simulation model, CTG++, which can generate instruction-compliant, realistic, and controllable traffic scenarios. Wang et al.[105] utilized natural language descriptions as conceptual representations that were integrated with LLM to enrich the complexity of the generated scenarios leveraging their powerful common-sense reasoning capabilities. The behavior of human drivers is also an important part of driving scenarios, Jin et al.[134] proposed SurrealDriver, a generative driving agent simulation framework for urban environments based on LLM. By analyzing and learning from real driving data, SurrealDriver can capture the driver’s behavioral patterns and decision-making processes and generate behavioral sequences that are similar to those in real driving. similar behavioral sequences as in real driving.
Through efficient and accurate simulation of driving scenarios, including a variety of complex and dangerous situations, a large amount of training data is provided for model learning, which can effectively enhance the generalization capability of autonomous driving systems. Meanwhile, the generated driving scenarios can also be used to evaluate different autonomous driving systems and algorithms to test and verify their system performance.
Table 2: Works on Data Augmentation
Authors Input Output Learning
Expand Dataset Qian et al.[115] 2023 Images, Text, Point Clouds Q-A Pair SL
Wu et al.[116] 2023 Images, Text Object-Text Pairs SL
Zhou et al.[117] 2023 Images, Text Labeled Data SL
Generate Scenarios Marathe et al.[129] 2023 Objects, Scenes, Weather Condition Multi-Weather Images SL
Chen et al.[132] 2023 Objects, Text Trajectory SL
Zhong et al.[133] 2023 Text, Scenes with Noise Traffic Scenes SL
Wang et al.[105] 2023 Images, Knowledge Latent Space Simulation SL
$\bullet$ This “SL” refers to “Supervised Learning” and “Q-A” refers to “Question-Answer”.
4 World Model
@@
World Models refer to the mental models of the world. It can be interpreted as a type of artificial intelligence model that encompasses a holistic understanding or representation of the environment in which it operates. This model is capable of simulating the environment to make predictions or decisions. The term “World Models” has been mentioned in the context of reinforcement learning in recent literature[137, levine2021understanding]. This concept has also gained attention in autonomous driving applications due to its capacity to comprehend and articulate the dynamics of the driving environment, as will be detailed below. In his position paper, LeCun[135] pointed out that the learning capability of humans and animals may be rooted in their capacity to learn World Models, allowing them to internalize and understand how the world works. He pointed out that humans and animals have exhibited the ability to acquire a vast amount of background knowledge about the functioning of the world observing a small number of events, whether related or unrelated to the task at hand.
The idea of World Model can be traced back as far as Dyna, proposed by Sutton et al.[136] in 1991, where observing the state of the world and selecting an action based on its response learns interactively with the world. Dyna is essentially a form of reinforcement learning under supervised conditions. After that, researchers have also made many attempts. Ha et al.[137] attempted to learn leveraging an unsupervised approach, VAE to encode input features and RNN to learn the evolution of the state. Hafner et al.[138] proposed the Recurrent state space model (RSSM), which combined reinforcement learning to realize multi-step prediction that integrates stochasticity and determinism. Based on the RSSM architecture, Hafner et al. successively proposed DreamerV1[139], DreamerV2[140], and DreamerV3[141], which learned in implicit variables to realize image prediction generation. Gao et al.[142] considered that there was redundant information implicit, and extended the framework of the Dreamer series by proposing the Semantic Masked recurrent World Model (SEM2) to learn relevant driving states. Hu et al.[143] removed prediction rewards and proposed an imitation learning-based method MILE to predict future states.
@@
It can be seen that World Model is highly related to reinforcement learning, imitation learning, and deep generative models. However, utilizing World Models in reinforcement learning and imitation learning generally requires labeled data, and both SEM2 and MILE approaches mentioned are conducted within a supervised paradigm. There have also been attempts to combine reinforcement learning and unsupervised learning based on the limitations of labeled data[144, 145]. Due to the close relationship with self-supervised learning, deep generative models have become more and more popular, and many attempts have been made by researchers in this field. In the following, we will mainly review the exploratory applications of generative World Models in autonomous driving, the pipeline is illustrated in Fig. 5, 4.1 introduces the principles of various types of deep generative models and their applications in generative driving scenarios, 4.2 introduces the applications of generative World Models in autonomous driving, and 4.3 will introduce a class of non-generative methods.
@@
Figure 5: The pipeline diagram for enhancing autonomous driving with World Models. The World Models first learn the intrinsic evolutionary patterns by observing the traffic environment and then enhance the autonomous driving by hooking up different decoders adapted to different driving tasks.
@@ 4.1 Deep Generative Models
Deep generative models generally include variational autocoders (VAEs) [146, 147], generative adversarial networks (GANs)[26, 148], flow models[149, 150], and autoregressive models (ARs)[8, 151, 152].
VAEs combine the ideas of self-encoders and probabilistic graphical models to learn the underlying structure of the data and generate new samples. Rempe et al.[153] used VAE to learn prior distributions of traffic scenarios and simulate the generation of accident-prone scenarios. GANs consist of two neural networks, the Generator and the Discriminator, which compete and enhance each other utilizing adversarial training, to ultimately achieve the goal of generating realistic samples. Kim et al.[154] used a GAN model to observe sequences of unlabeled video frames and their associated action pairs to simulate a dynamic traffic environment. The flow models generate similar data samples by transforming simple prior distributions into complex posterior distributions through a series of invertible transformations. Kumar et al.[155] used the flow model to achieve multi-frame video prediction. ARs are a class of sequence analysis methods, based on the autocorrelation between the sequence data, describing the relationship between the current observations and the past observations, and the estimation of the model parameters is usually done leveraging the least squares method and maximum likelihood estimation. For example, GPT uses maximum likelihood estimation for model parameter training. Feng et al.[156] achieved the generation of future trajectories of vehicles based on autoregressive iterations. The diffusion Model is a typical autoregressive model that learns the process of gradual denoising from purely noisy data. With its strong generative performance, the Diffusion Model is the new SOTA among current deep generative models. Works such as[157, 158, 159] demonstrated that the Diffusion Model has a strong ability to understand complex scenes, and the video diffusion model can generate higher quality videos. Work such as[160, 161] utilized the Diffusion Model to generate complex and diverse driving scenarios.
@@ 4.2 Generative Methods
Based on the powerful capabilities of deep generative models, leveraging deep generative models as World Models to learn driving scenarios to enhance autonomous driving has become popular. The following section will review the applications of leveraging deep generative models as World Models in autonomous driving. In Table LABEL:tab:3 we provide a brief overview of some representative works.
@@
Karlsson et al.[162] used a hierarchical VAE to construct World Model, used latent variable prediction and adversarial modeling to generate pseudo-complete states, matched partial observations with pseudo-complete observations to predict future states, and evaluated it on the KITTI-360[163] dataset. In particular, it utilized pre-trained vision-based semantic segmentation models to infer semantic information from image observations without manual annotation. Similarly, to address the challenges of predicting future changes in driving scenarios, Wayve proposed a generative World Model, GAIA-1[7]. GAIA-1 used transformer as World Model, learned and predicted the next states of the input video, text, and action signals and then generated a realistic driving scenarios. For the learning of video streams, GAIA-1 adopted self-supervised learning, which can learn scaled data and obtain a comprehensive understanding of driving scenarios. Wang et al[164] designed a two-phase training strategy that incorporates the features of the diffusion model to learn the driving scenarios and obtain an understanding of the structured traffic information. The driving World Model is then modeled through video prediction task, specifically, the prediction is formalized into a generative probabilistic model to generate future driving scenarios. Moreover, by integrating historical driving behavior, this method can also generate future driving actions. zhang et al.[165] built on Maskgit[166] and recast it into a discrete diffusion model for point clouds prediction. This method utilized VQ-VAE[167] to tokenize the observation data for label-free learning. Bogdoll et al[bogdoll2023muvo] constructed a multimodal autonomous generative World Model, MUVO, leveraging raw images and LiDAR data to learn a geometric representation of the world. And conditioned on actions, this mode achieved 3D occupancy prediction and can be directly used for downstream tasks(e.g.,planning). Similarly, Min et al.[min2023uniworld] used unlabeled image-LiDAR for pre-training to construct World Model that can generate 4D geometric occupancy. This work intends to obtain a FM through pre-training and subsequently fine-tuning on downstream tasks. In the industry, at the 2023 CVPR Autonomous Driving Workshop, Tesla researcher Ashok presented exploratory work in utilizing generative large model[168] to generate future driving scenarios. And in the demonstration it was seen that the videos generated by Tesla’s generative large model were very close to those captured from real vehicles. It also can generate annotation-like semantic information, indicating that the model also has some semantic-level understanding and reasoning capabilities. Tesla named their work “Learning a General World Model” and it can be seen that their understanding is to build a generalized World Model. By learning from a large amount of video data captured from real vehicles, Tesla intends to build a large-scale FM for autonomous driving, which can understand the dynamic evolution of the world by learning from image data at scale.
@@
Vision is one of the most direct and effective ways for humans to obtain information about the world because the feature information contained in image data is extremely rich. Numerous previous works[139, 140, 141, 145, 169] have accomplished the task of image generation through World Model, demonstrating that World Model has a good understanding and reasoning ability for image data. However, these are mainly focused on image generation and are still lacking in video prediction tasks that can better represent the dynamic evolution of the world. Video prediction tasks require a deeper understanding of world evolution and also stronger guidance for downstream tasks. In the research works[7, 162], they all effectively predicted generated future traffic scenarios, where self-supervised learning may be key. Previous work has explored this as well. Villegas et al.[170] trained a model leveraging raw images and proposed a hierarchical long-term video prediction method combining low-level pixel space and high-level feature space (e.g., landmarks), achieving longer video prediction compared to the work[141]. Endo et al.[171] constructed a model under the self-supervised learning paradigm for predicting future traffic scenes from single-frame images to predict the future. Based on a denoising diffusion model with probabilistic conditional scores, Voleti et al.[172] trained the model by randomly masking the past frames or future frames unlabeled, which allowed block-by-block autoregressions to generate videos of arbitrary length. Finn et al.[173] proposed to physically interact with the world under unsupervised conditions and realized video prediction by predicting the distribution of pixel motion in the previous frame. Micheli et al.[174] verified the effectiveness of leveraging an autoregressive transformer as a World Model, and achieved the prediction of game images by training the parameters through self-supervised learning. Wu et al.[175] constructed an object-centered World Model to learn complex spatio-temporal interactions between objects and generated high visual quality future prediction.
From the above research, we can see that researchers hope to learn the intrinsic evolutionary laws of the world from image data, and then predict the future state. Combined with self-supervised learning, the World Model is used to learn from image data, fully unleashing the inference ability of the model, and constructing a generalised Foundation Model for the visual domain may be promising.
@@
AD Wang et al.[164] 2023 Images, HDMap, 3D Box, Text, Actions Videos, Actions SSL This work obtains comprehension of the structured traffic information. Then the prediction is formalized into a generative probabilistic model.

### v2

@@ 3.3 Generation of Actions
Xu et al.[75] 2023 Video, Text Prompt, Control Signal Decision, Control Signal SL This work tokenizes video sequences, text, and control signals to build the model, which can generate responses to human inquiries and predict control signals.
Sima et al.[drivelm] 2023 Video, Text Question Scene Description, Decision, Trajectory SL Based on Graph Visual Question Answering (GVQA), this work realizes structured reasoning for perception, prediction, and planning through suitable quizzes for human-like autonomous driving.
Shao et al.[lmdrive] 2023 Camera, LiDAR, Text Prompt Trajectory SL This work accomplishes e2e autonomous driving by interacting with dynamic environments through multi-modal multi-view sensor data and language commands.
Ma et al.[ma2023dolphins] 2023 Video, Text Prompt, Control Signal Scene Description, Prediction, Trajectory SL This work employs a Grounde-CoT process to enhance the model’s reasoning capabilities. This work also integrates four different tasks to facilitate the model’s comprehensive understanding of complex driving scenarios.
4 Prediction of Autonomous Driving Based on World Models
@@
World Models refer to the mental models of the world. It can be interpreted as a type of artificial intelligence model that encompasses a holistic understanding or representation of the environment in which it operates. This model is capable of simulating the environment to make predictions or decisions. The term “World Models” has been mentioned in connection with reinforcement learning in recent literature[137, levine2021understanding]. This concept has also gained attention in autonomous driving because of its capacity to comprehend and articulate the dynamics of the driving environment, as will be detailed below. In his position paper, LeCun[135] pointed out that the learning capability of humans and animals may be rooted in their capacity to learn World Models, allowing them to internalize and understand how the world works. He pointed out that humans and animals have exhibited the ability to acquire a vast amount of background knowledge about the functioning of the world observing a small number of events, whether related or unrelated to the task at hand. The idea of World Model can be traced back to Dyna, proposed by Sutton et al.[136] in 1991, to observe the state of the world and take appropriate actions accordingly to learn interactively with the world[fan2023advanced]. Dyna is essentially a form of reinforcement learning under supervised conditions. After that, researchers have also made many attempts. Ha et al.[137] attempted to learn by leveraging an unsupervised approach, VAE to encode input features, and RNN to learn the evolution of the state. Hafner et al.[138] proposed the Recurrent State Space Model (RSSM), which combined reinforcement learning to realize multi-step prediction that integrates stochasticity and determinism. Based on the RSSM architecture, Hafner et al. successively proposed DreamerV1[139], DreamerV2[140], and DreamerV3[141], which learned in implicit variables to realize image prediction generation. Gao et al.[142] considered that there was redundant information implicit, and extended the framework of the Dreamer series by proposing the Semantic Masked recurrent World Model (SEM2) to learn relevant driving states. Hu et al.[143] removed prediction rewards and proposed an imitation learning-based method MILE to predict future states.
@@
It can be seen that World Model is highly related to reinforcement learning, imitation learning, and deep generative models. However, utilizing World Models in reinforcement learning and imitation learning generally requires labeled data, and both SEM2 and MILE approaches mentioned are conducted within a supervised paradigm. There have also been attempts to combine reinforcement learning and unsupervised learning (UL) based on the limitations of labeled data[144, 145]. Due to the close relationship with self-supervised learning, deep generative models have become more and more popular, and researchers in this field have made many attempts. In the following, we will mainly review the exploratory applications of generative World Models in autonomous driving, the pipeline is illustrated in Fig. 7, 4.1 introduces the principles of various types of deep generative models and their applications in generative driving scenarios, 4.2 introduces the applications of generative World Models in autonomous driving, and 4.3 will introduce a class of non-generative methods.
@@
Figure 7: The pipeline diagram for enhancing autonomous driving with World Models. The World Models first learn the intrinsic evolutionary patterns by observing the traffic environment and then enhance autonomous driving by hooking up different decoders adapted to different driving tasks.
@@ 4.1 Deep Generative Models
Deep generative models generally include ariational autocoders (VAEs) [146, 147], generative adversarial networks (GANs)[26, 148], flow models[149, 150], and autoregressive models (ARs)[8, 151, 152].
VAEs combine the ideas of self-encoders and probabilistic graphical models to learn underlying data structures and generate new samples. Rempe et al.[153] used VAE to learn prior distributions of traffic scenarios and simulate the generation of accident-prone scenarios. GANs consist of generator and discriminator, which compete and enhance each other utilizing adversarial training, to ultimately achieve the goal of generating realistic samples. Kim et al.[154] used a GAN model to observe sequences of unlabeled video frames and their associated action pairs to simulate a dynamic traffic environment. The flow models generate similar data samples by transforming simple prior distributions into complex posterior distributions through a series of invertible transformations. Kumar et al.[155] used the flow model to achieve multi-frame video prediction. ARs are a class of sequence analysis methods, based on the autocorrelation between the sequence data, describing the relationship between the present and the past, and the estimation of the model parameters is usually done leveraging the least squares method and maximum likelihood estimation. For example, GPT uses maximum likelihood estimation for model parameter training. Feng et al.[156] achieved the generation of future trajectories of vehicles based on autoregressive iterations. The diffusion model is a typical autoregressive method that learns the process of gradual denoising from purely noisy data. With its strong generative performance, the diffusion model is the new SOTA among current deep generative models. Works such as[157, 158, 159] demonstrated that the diffusion model has a strong ability to understand complex Scenarios, and the video diffusion model can generate higher quality videos. Work such as[160, 161] utilized the diffusion model to generate complex and diverse driving scenarios.
@@ 4.2 Generative Methods
Based on the powerful capabilities of deep generative models, leveraging deep generative models as World Models to learn driving scenarios to enhance autonomous driving has become popular. The following section will review the applications of leveraging deep generative models as World Models in autonomous driving. In Table LABEL:tab:2 we provide a brief overview of some representative works.
@@
Point Cloud involved Models. Zhang et al.[165] built on Maskgit[166] and recast it into a discrete diffusion model for point cloud prediction. This method utilized VQ-VAE[167] to tokenize the observation data for label-free learning. Karlsson et al.[162] used a hierarchical VAE to construct World Model, used latent variable prediction and adversarial modeling to generate pseudo-complete states, matched partial observations with pseudo-complete observations to predict future states, and evaluated it on the KITTI-360[163] dataset. In particular, it utilized pre-trained vision-based semantic segmentation models to infer from raw images. Bogdoll et al[bogdoll2023muvo] constructed a multi-modal autonomous generative World Model, MUVO, leveraging raw images and LiDAR data to learn a geometric representation of the world. And conditioned on actions, this mode achieved 3D occupancy prediction and can be directly applied to downstream tasks (e.g., planning). Similarly, Zheng et al.[zheng2023occworld] used VQ-VAE to tokenize the 3D occupancy scene and constructed a 3D occupancy apace to learn a World Model that can predict the motion of the ego-vehicle and the evolution of the driving scenario. To obtain finer-grained scene information, Min et al.[min2023uniworld] used unlabeled image-LiDAR for pre-training to construct World Model that can generate 4D geometric occupancy.
Image-based Models. To address the challenges of predicting future changes in driving scenarios, Wayve proposed a generative World Model, GAIA-1[7]. GAIA-1 used transformer as World Model learned and predicted the next states of the input video, text, and action signals, and then generated realistic driving scenarios. For the learning of video streams, GAIA-1 adopted self-supervised learning, which can learn scaled data and obtain a comprehensive understanding of driving scenarios. Wang et al.[164] devised a two-stage training strategy. Initially, a diffusion model was employed to learn driving scenarios and gain an understanding of structured traffic information. Subsequently, a video prediction task was used to construct a world model, designated DriveDreamer. Notably, by integrating historical driving behaviors, this approach enables the generation of future driving actions. Zhao et al.[zhao2024drivedreamer] constructed DriveDreamer-2 on top of the DriveDreamer framework by integrating LLM, which generates the corresponding agent trajectories based on user descriptions, and HDMap information to controllably generate driving videos. Wang et al.[wang2023driving] generated the driving videos by jointly modeling the future multi-views and multi-frames. This approach greatly improved the consistency of the generated results, and end-to-end motion planning was generated based on this.
In the industry, at the 2023 CVPR Autonomous Driving Workshop, Tesla researcher Ashok Elluswamy presented their work in utilizing generative large model[168] to generate future driving scenarios. In the demonstration, it was seen that the videos generated by Tesla’s generative large model were very close to those captured from real vehicles. It also can generate annotation-like semantic information, indicating that the model also has some semantic-level understanding and reasoning capabilities. Tesla named their work “Learning a General World Model” and it can be seen that their understanding is to build a generalized World Model. By learning from a large amount of visual data captured from real vehicles, Tesla intends to build a large-scale FM for autonomous driving, which can understand the dynamic evolution of the world.
@@
Visual Prediction. Vision is one of the most direct and effective means by which humans acquire information about the world because the feature information contained in image data is extremely rich. Numerous previous works[139, 140, 141, 145, 169] have accomplished the task of image generation through World Model, demonstrating that World Model has a good understanding and reasoning ability for image data. However, these are mainly focused on image generation and are still lacking in video prediction tasks that can better represent the dynamic evolution of the world. Video prediction tasks require a deeper understanding of world evolution and also stronger guidance for downstream tasks. In the research works[7, 162], they all effectively predicted generated future traffic scenarios, where self-supervised learning may be key. Previous work has explored this as well. Villegas et al.[170] trained a model leveraging raw images and proposed a hierarchical long-term video prediction method combining low-level pixel space and high-level feature space (e.g., landmarks), achieving longer video prediction compared to the work[141]. Endo et al.[171] constructed a model under the self-supervised learning paradigm for predicting future traffic Scenarios from single-frame images to predict the future. Based on a denoising diffusion model with probabilistic conditional scores, Voleti et al.[172] trained the model by randomly masking the past frames or future frames unlabeled, which allowed block-by-block autoregressions to generate videos of arbitrary length. Finn et al.[173] proposed to physically interact with the world under unsupervised conditions and realize video prediction by predicting the distribution of pixel motion. Micheli et al.[174] verified the effectiveness of leveraging an autoregressive transformer as a World Model, and achieved the prediction of game images by training the parameters through self-supervised learning. Wu et al.[175] constructed an object-centered World Model to learn complex spatio-temporal interactions between objects and generated high visual quality future prediction.
Inspired by LLM, Wang et al.[wang2024worlddreamer] consider world modeling as unsupervised visual sequence modeling. The visual input is mapped into discrete tokens using VQ-GAN[esser2021taming], and then the Spatio-Temporal Transformer is used to predict the masked tokens to learn the physical evolutionary patterns in them, thus gaining the ability to generate videos in various scenarios. Analogous to LLM’s tokens, OpenAI researchers transformed visual data into patches to propose the video generation model Sora. To address the high-dimensionality of visual data, they compressed the visual data into a lower-dimensional latent space and then generated a latent representation in this latent space through diffusion. This representation was then mapped back to the pixel space to realize video generation. By learning from Internet-scale data, Sora realizes the scaling law in the video domain, and Sora can generate coherent high-definition videos based on diverse prompts. In the same year, Google proposed Genie[bruce2024genie], a generative interactive model that uses unlabeled Internet gaming videos for training. In particular, Genie proposed a latent action model to infer latent actions between each frame and constructed a codebook for latent actions through training. To utilize the model, the user selects the initial frame and the specified latent action and autoregressively generates future frames. As the model size and batch size increase, Genie also demonstrates scaling results. In contrast, Sora is designed to generate video content with high fidelity, variable duration, and resolution. While not as advanced in video quality as Sora, Genie is optimized for building generative interactive environments in which the user can manipulate frame-by-frame to generate video.
The preceding studies demonstrate the efficacy of World Models in enhancing autonomous driving. World Models can be directly embedded into autonomous driving models to accomplish various driving tasks. Furthermore, there are explorations of learning to build general World Models from large-scale visual data, such as Sora and Genie. These FMs can be utilized for data generation (to be discussed in Section 5). In addition, based on FMs’ generalization ability, they can be employed to perform a multitude of downstream tasks, or even be utilized to simulate the world.
@@
Wang et al.[164] 2023 Images, HDMap, 3D Box, Text, Actions Videos, Actions SSL This work obtains comprehension of the structured traffic information. Then the prediction is formalized into a generative probabilistic model.